OpenStudy (anonymous):

$6,000 for 6 years at 8 1/2% compounded daily will grow to

OpenStudy (anonymous):

\[A = P(1 + \frac{r}{n})^{nt}\]\[A = 6000(1 + \frac{0.085}{365})^{365(6)}\]

OpenStudy (anonymous):

Can you solve form here?